My teacher asked me this question
3x + 5y = 4
2x + 4y = 6
She said I could use either substitution or elimination. I want to show both, can you help me solve this equation using both substitution and elimination (separately) ?

Answer by Boreal(15235)

3x + 5y = 4
2x + 4y = 6
=========
6x+10y=8
-6x-12y=-18
-2y=-10
y=5
2x+20=6
2x=-14
x=-7
(-7.5)
Check with substitution
====================
2x+4y=6
2x=-4y+6
x=-2y+3
substitute into first equation
3(-2y+3)+5y=4
-6y+9+5y=4
-y=-5
y=5
x will be -7.